OpenStudy (anonymous):

Sally made the conjecture that the sum of any two prime numbers is a prime number. Which equation is a counterexample to her conjecture? A. 15 + 4 = 19 B. 2 + 11 = 13 C. 3 + 11 = 14 D. 2 + 5 = 7

OpenStudy (anonymous):

@Ghhhostt

OpenStudy (anonymous):

can anyof yall help

OpenStudy (anonymous):

Which one of the answers is not a prime number?

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A and D

OpenStudy (matlee):

prime numbers are numbers that i think cannot be divided by another number other than one so just look through those

OpenStudy (anonymous):

A and D

OpenStudy (anonymous):

Prime numbers are numbers that are divisible by the number itself and the number 1. Prime numbers go on forever, but here is a list of prime numbers up to 20: 2, 3, 5, 7,11,13,17,19 So we are looking for an answer that makes the conjecture not true. Looking at the first answer, we see that when we add two prime numbers we get a number that is not on our list of prime numbers. So this answer shows that the conjecture is not true. This answer is the correct answer. Looking at the other answers, we see that when we add two prime numbers we get a prime number. So this answer doesn't show the conjecture is not true. So this answer is no good. The third answer again shows that when we add two prime numbers we get another prime number. So this answer doesn't show the conjecture is not true. So this answer is no good. The final answer adds a prime number and a composite number. But the conjecture says we must add two prime numbers. So this answer is no good. So the answer C is correct.

OpenStudy (anonymous):

csomeone medal @Ghhhostt and @matlee

OpenStudy (matlee):

doesnt the question say that all number have to be prime?

OpenStudy (matlee):

i thought it was D

OpenStudy (anonymous):

A because a counter example is something that is the opposite and 15 is not a prime number

OpenStudy (matlee):

oh yeah counter example

OpenStudy (anonymous):

Lol im glad you all are so quick >.< My website keeps freezing randomly.

OpenStudy (anonymous):

so i am confused A C or D

OpenStudy (anonymous):

I would say C

OpenStudy (matlee):

i agree with A

OpenStudy (matlee):

lol this is more complicated than i thought

OpenStudy (anonymous):

Sally made the conjecture that the sum of any two prime numbers is a prime number. Which equation is a counterexample to her conjecture? A. 15 + 4 = 19 B. 2 + 11 = 13 C. 3 + 11 = 14 D. 2 + 5 = 7

OpenStudy (anonymous):

A or C

OpenStudy (matlee):

Ayyyyyy

OpenStudy (anonymous):

The problem with A being an example is that it's not a counter. 4 is not a prime number so it's cannot fit. I say C

OpenStudy (matlee):

nvm it is C

OpenStudy (matlee):

Definition of COUNTEREXAMPLE. : an example that refutes or disproves a proposition or theory.

OpenStudy (anonymous):

matlee aren't you a human calculator

OpenStudy (anonymous):

^_^

OpenStudy (matlee):

Yes but these are words!

OpenStudy (matlee):

:(

OpenStudy (anonymous):

Yep. Its C for sure.

OpenStudy (anonymous):

ahahha he's got a point

OpenStudy (matlee):

-pulls out glossary for math class

OpenStudy (anonymous):

you can use a calculator

OpenStudy (matlee):

The answer is C

OpenStudy (anonymous):

Yes, it's C for sure. They're trying to be sneaky and put that 4 into the eq.

OpenStudy (anonymous):

Knew it all along :D

OpenStudy (anonymous):

C it is!

OpenStudy (anonymous):

ok stay so i can tell if it was right

OpenStudy (matlee):

Ok

OpenStudy (anonymous):

AWESOME you were right C

OpenStudy (anonymous):

KNEW IT

OpenStudy (anonymous):

Yay I happy 4 you. Good work! :D

OpenStudy (anonymous):

thanks

OpenStudy (matlee):

Good job Denisha

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Lol I was about to ask how you know my name and then I looked up at my user xD